OpenStudy (anonymous):

I have 2 questions :) Appreciate any help Tell whether the lines for each pair of equations are parallel perpendicular or neither. y=1/4x+10 -2x+8y=6 a parallel b. perpendicular c. neither

OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line that passes through the given point. y-3=8/3(x+2);(-2,3) a. y+3=-3/8(x-2) b. y-3=-3/8(x+2) c. y-2=-3/8(x+3) d. y+3=3/8(x-2)

OpenStudy (anonymous):

Will fan and medal

OpenStudy (sam141101):

sorry I'm not really good at this sorry

OpenStudy (anonymous):

Okay

OpenStudy (sam141101):

sorry

OpenStudy (anonymous):

I'm not either,...at all lol

OpenStudy (sam141101):

really same

OpenStudy (sam141101):

im not good at math at all